Find an equation for the line with y-intercept 3 that is perpendicular to the line y=(2/3)x-4. Thank you!
Answer by jake_6233(74)

Find an equation for the line with y-intercept 3 that is perpendicular to the line y=(2/3)x-4.
y=mx+b is important to remember b is the y intercept and m is the slope
m*m1=-1 is also important here for the slope of the requested line you want the negative reciprocal (flip the fraction and change the sign)
y= -(3/2)x -3 I hope this helps